OpenStudy (anonymous):

how many solutions does this system have x-2y=2 y=-2+5 A.infinitely many solutions B.no solution C.two solutions B.one solution

OpenStudy (freckles):

is the second equation really y=3 ?

OpenStudy (anonymous):

x-2y=2 y=-2x+5 I messed up

OpenStudy (freckles):

since you don't actually need to find the solution(s) I would just put both in slope-intercept form and compare slopes then intercepts if necessary what I mean is you can have the following possibilities: A) Different slopes means one solution B) Same slopes and same y-intercept means infinite amount of solutions C) Same slopes and different y-intercepts means no solution

OpenStudy (freckles):

so y=-2x+5 is already in slope-intercept form

OpenStudy (freckles):

x-2y=2 is not

OpenStudy (freckles):

to get x-2y=2 in slope intercept form we want to write in y=mx+b form that means we will need to subtract x on both sides then divide both sides by -2 x-2y=2 subtract x on both sides -2y=-x+2 divide both sides by -2 y=1/2*x-1

OpenStudy (freckles):

now what is the slope of y=-2x+5 and what is the slope of y=1/2*x-1

OpenStudy (anonymous):

the first one is -2 the second one is 1/2 @freckles

OpenStudy (freckles):

and so if they have different slopes how many solutions with the system of linear equations have?

OpenStudy (freckles):

refer to A) Different slopes means one solution B) Same slopes and same y-intercept means infinite amount of solutions C) Same slopes and different y-intercepts means no solution

OpenStudy (anonymous):

one solution

OpenStudy (freckles):

yep yep

OpenStudy (anonymous):

Thanks